Fermi National Laboratory


Accelerator Update for December 15, 2004 - December 17, 2004

Wednesday December 15
The day shift began with Operations monitoring the Antiproton Source (Pbar) stack, with the Main Injector (MI) conducting no-ramp studies, and with the Booster, Tevatron (TeV), and Switchyard (SY) in access for maintenance work.

Operators prepared MI for access at 9:15 AM.

Operations established beam first to the Booster dump and then to MiniBooNE at 12:45 PM.

Operations established beam to MI and then to Pbar, which resumed stacking at 1 PM.

The Recycler took protons for studies at 2 PM.

At 3:28 PM, Operations reported that problems with a MI power supply (I:SEXVPS). Experts investigated and reestablished beam in MI at 5 PM. Pbar resumed stacking.

The Collider Detector at Fermilab (CDF) and D-Zero (D0) experimenters were out of access by 5:40 PM.

Operations put the Tevatron (TeV) through a dry squeeze at 7:47 PM.

Operations put the TeV into shot setup at 9:36 PM.

Operations established store 3850 at 11:59 PM with an initial luminosity of 51.3E30.

Thursday December 16
The midnight shift began with Operations monitoring store 3850, with Pbar returning to stacking, and with MiniBooNE taking beam.

At 4:45 AM, Operations reported a spike in the colliding beam that both experiments saw. Operator investigated and found a possible problem with the sector A4 separator; Whatever it was luminosity wasn't affected.

At 2:05 PM, the TeV suffered a quench at sector F4 due to a B1 separator spark.

Operators made D0 ready for an access at 2:54 PM.

A NuMI expert began a study period at 3:17 PM.

At 4:05 PM, the MI ramp tripped off due to a MECAR problem.

MI beam returned and Pbar resumed stacking at 4:30 PM.

Operations reported at 6:18 PM that the TeV was being held off by a mechanical dump switch problem and a separator power supply problem.

A TeV expert began an orbit study at 9:27 PM.

Operations put the TeV into shot setup at 10:34 PM.

At 11:32 PM, Operations reported that problems with a TeV reverse proton tune up had put shot setup on hold.

Friday December 17
The midnight shift began with the TeV in shot setup and with Pbar having a turn-by-turn problem.

The Pbar problem was resolved by 1:11 AM.

Operations established store 3853 at 2:11 AM with an initial luminosity of 54.6E30. (This is the highest luminosity since the shutdown.) Pbar resumed stacking.

After resolving some problems, MI began sending beam to SY at 3:15 AM.

At 2:56 AM, Operations reported the first of what seem to be many A4 separator sparks, but there was no luminosity loss.

Plans for Friday and the Weekend
The plan for today is to keep the store in until this evening and let Pbar build up a big stack, then put the TeV into shot setup. The plans for the weekend are to stack and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/17/2004   email Fermilab